OpenStudy (anonymous):

help...Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1 fourthx + 2. The answer choices are y = 4x -1 y = 4x - 3 y = 1/4x - 3 y = 1/4x - 1

OpenStudy (anonymous):

I know it can't be A or B because they do not have the same slope as the original equation and just saying the slope is 1/4

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

Good observation... now select equation C or D and plug in the values (-4, -2) and see which one C or D gives you a true equation.. let me know if you get it, this is my first time on this site and I am still trying to figure out how to work the site

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

how would i plug in the points

OpenStudy (anonymous):

help

OpenStudy (anonymous):

if they gave you points (x,y) plug in coordinate y for y in the equation and coordinate x for x in the equation

jimthompson5910 (jim_thompson5910):

Anything parallel to this line will have a slope of 1/4 (since parallel lines have equal slopes) So m = 1/4 The given point is (-4,-2). So (x,y) = (-4,-2) which means x = -4 and y =-2 Plug all this into y = mx+b and solve for b y = mx+b -2 = (1/4)(-4)+b -2 = (1/4)(-4)+b -2 = (1/4)(-4/1)+b -2 = (1*(-4))/(4*1)+b -2 = (-4)/(4)+b -2 = -1+b -2 + 1 = b -1 = b b = -1 ------------------------- So we know that m = 1/4 and b = -1 That means the final answer is \(\Large y = \frac{1}{4}x - 1\)

OpenStudy (anonymous):

omg thanks you get a medal

jimthompson5910 (jim_thompson5910):

you're welcome, and thanks